If you cut out the largest possible circular pie crust from a square piece of dough that is d inches on a side, what is the area of the remaining dough?
If you take this remaining dough and reshape it into a new circle to make a second pie, what is the diameter of the second pie?

Answer by checkley71(8403)

D^2=THE AREA OF THE SQUARE.
D^2-PI(D/2)^2 IS THE AMOUNT LEFT OVER AFTER CUTTING OUT THE CIRCLE OR THE AREA OF THE NEW CIRCLE.
[D^2-PI(D/2)^2]=PI(d/2)^2 {d being the diameter of the smaller left over pie}
[D^2-PI(D/2)^2]/PI=(d/2)^2
d/2=SQRT{[D^2-PI(D/2)^2]/PI}
d=2SQRT{[D^2-PI(D/2)^2]/PI} ANSWER FOR THE DIAMETER OF THE SECOND PIE.